# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ Appendix C Detailed Proofs for the Policy Gap Bound (Theorem 1) and Supporting Results
Throughout the proof, we abbreviate $\lVert\cdot\rVert_{q,p}:=\lVert\cdot\rVert_{L^{q,p}}$. For CRRA utility $U(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, we have $U^{\prime}(c)=c^{-\gamma}$ and $(U^{\prime})^{-1}(\ell)=\ell^{-1/\gamma}$. By (31), $\sup_{c\in[\underline{c},\overline{c}]}\!U^{\prime}(c)=\underline{c}^{-\gamma}\;<\;\infty$, and $\sup_{\ell\in[\underline{\lambda},\overline{\lambda}]}\bigl{|}(U^{\prime})^{-1}\!{}^{\prime}(\ell)\bigr{|}=\frac{1}{\gamma}\,\underline{c}^{\gamma}\,\overline{\lambda}^{\frac{\gamma+1}{\gamma}}=:L_{(U^{\prime})^{-1}}$. Hence $(U^{\prime})^{-1}$ is globally Lipschitz on $[\underline{\lambda},\overline{\lambda}]$ with constant $L_{(U^{\prime})^{-1}}$.
Next, by Taylor expansion of $U$, for every $(t,\omega)$ there exists $c^{\dagger}_{t,\omega}\in[\underline{c},\overline{c}]$ such that $U(c^{*}_{t})-U(c_{t})=U^{\prime}(c^{\dagger}_{t,\omega})\,(c^{*}_{t}-c_{t})$, and $\bigl{|}U^{\prime}(c^{\dagger}_{t,\omega})\bigr{|}\leq\underline{c}^{-\gamma}$. Multiplying by $e^{-\delta t}$ and taking absolute values yields $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\underline{c}^{-\gamma}\,e^{-\delta t}\,\bigl{|}\,c^{*}_{t}-c_{t}\bigr{|}$.
Furthermore, the first-order condition for consumption gives $e^{-\delta t}U^{\prime}(c^{*}_{t})=\lambda^{*}_{t}$ and $e^{-\delta t}U^{\prime}(c_{t})=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$, where $\mathcal{R}_{\mathrm{FOC}}^{(c)}$ denotes the consumption component of the FOC–residual. Because both arguments $\ell^{*}:=\lambda^{*}_{t}$ and $\ell:=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$ lie in $[\underline{\lambda},\overline{\lambda}]$, the Lipschitz property of $(U^{\prime})^{-1}$ yields $|c^{*}_{t}-c_{t}|\;\leq\;L_{(U^{\prime})^{-1}}\,\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$.
Combining these results, we insert the bound for $|c^{*}_{t}-c_{t}|$ into the inequality for $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}$ to get $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\Bigl{(}\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\Bigr{)}\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$. Taking the $L^{q,p}$–norm, using the definitions of $E_{v}$ and $\varepsilon$, and noting $|\lambda-\lambda^{*}|=|V^{\pi,c}_{x}-V^{*}_{x}|\leq|\nabla V^{\pi,c}-\nabla V^{*}|$, we conclude $E_{f}\;\leq\;\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\,\varepsilon\;+\;L_{(U^{\prime})^{-1}}\,E_{v}$. With $C_{f},C_{f}^{\prime}$ chosen as in the statement, (32) follows. ∎
Proposition C.1 (Value– and gradient–gap stability, full version).
Fix exponents $p>d+2,\;q\geq 2$. Under Assumption 6 (coefficient regularity, uniform parabolicity) and with the notation
$E_{f}:=\|f\|_{L^{q,p}},\qquad E_{h}:=\|h\|_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
one has
$E_{v}\;\leq\;C_{S}\,\bigl{(}E_{f}+E_{h}\bigr{)},$ (34)
where the stability constant $C_{S}$ depends only on $p,q,T,d,\underline{\nu},\|a^{*}\|_{L^{\infty}},\|b^{*}\|_{L^{\infty}}$ ($\underline{\nu}$: the uniform ellipticity constant; $a^{*},b^{*}$: diffusion and drift coefficients of $\mathcal{L}_{\pi^{*},c^{*}}$).
@@
First, we consider the equation for the difference $W:=V^{\pi,c}-V^{*}$. Recall that, for each admissible control, $V^{\pi,c}$ solves the linear backward parabolic PDE $-\partial_{t}V^{\pi,c}+{\mathcal{L}}_{\pi,c}V^{\pi,c}=f_{\pi,c}$, with $V^{\pi,c}(T,\cdot)=g$, where the terminal datum is $g(x)=\kappa U(x)$. Subtracting the optimal‐control equation yields $-\partial_{t}W+{\mathcal{L}}_{\!*}W=f+h\cdot\nabla V^{\pi,c}$, with $W(T,\cdot)=0$. Here $\mathcal{L}_{\!*}:=\mathcal{L}_{\pi^{*},c^{*}}$, $f:=f_{\pi,c}-f_{\pi^{*},c^{*}}$, and $h:=(b_{\pi,c}-b_{\!*},a_{\pi,c}-a_{\!*})$ abbreviates the vector of coefficient gaps; thus $h\in L^{q,p}$ by Lemma 2.
Regarding the coefficient class, Assumption 6(a) implies that $(t,x)\mapsto a_{\!*}(t,x)$ is uniformly VMO in $x$ with bounded measurable $t$–dependence, and satisfies the ellipticity bound $\underline{\nu}\,|\,\xi|^{2}\leq\xi^{\!\top}a_{\!*}(t,x)\xi\leq\|a_{\!*}\|_{\infty}|\,\xi|^{2}$. The lower order coefficients $b_{\!*}$ are bounded.
For the source term, define $g_{src}:=f+h\cdot\nabla V^{\pi,c}$ (using $g_{src}$ to avoid conflict with terminal data $g$). By Lemma 3 and Lemma 2, $f,h\in L^{q,p}$. Moreover, $\nabla V^{\pi,c}\in L^{q,p}$ (by maximal‐regularity for $V^{\pi,c}$), hence $g_{src}\in L^{q,p}$ and $\lVert g_{src}\rVert_{q,p}\leq E_{f}+E_{h}$.
Then, we obtain an $L^{q,p}$–gradient estimate. The equation for $W$ is a linear, non–divergence form, backward parabolic PDE with VMO diffusion coefficients, bounded drift, zero terminal data, and right–hand side $g_{src}\in L^{q,p}$. Therefore, Dong–Kim (Dong and Kim, , 2009, Thm. 2.1) (adapted to the backward direction by time reversal) yields $\|\nabla W\|_{L^{q,p}}\;\leq\;C_{\!DK}\,\|g_{src}\|_{L^{q,p}}$, where $C_{\!DK}=C_{\!DK}(p,q,d,T,\underline{\nu},\|a_{\!*}\|_{\infty},\|b_{\!*}\|_{\infty})$.
Finally, inserting the bound on $g_{src}$ and setting $C_{S}:=C_{\!DK}$ leads to the conclusion (34). ∎
Proposition C.2 (Coupled error‐estimate resolution, full version).
Let the constants $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$ be those obtained in Lemmas 3, 2 and Proposition C.1. Define $\rho:=C_{S}\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$. Then, under Assumptions 67, the error triple $(E_{f},E_{h},E_{v})$ satisfies
$E_{f}\;\leq\;C_{f}^{\prime\prime}\varepsilon,\quad E_{h}\;\leq\;C_{h}^{\prime\prime}\varepsilon,\quad E_{v}\;\leq\;C_{v}^{\prime\prime}\varepsilon,$ (35)
with finite constants
$C_{v}^{\prime\prime}=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\!*}},\quad C_{f}^{\prime\prime}=C_{f}+C_{f}^{\prime}C_{v}^{\prime\prime},\quad C_{h}^{\prime\prime}=C_{\mathrm{FOC}}+L_{H}C_{v}^{\prime\prime},$
where if $\rho<1$, we may take $\rho_{\!*}:=\rho$; otherwise, if $\rho\geq 1$, pick an integer $m\geq 1$ such that $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}\;<1$ for $\Delta:=T/m$ (possible because $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$). Then set $\rho_{\!*}:=\rho_{\Delta}$ and $C_{S}:=\max_{0\leq k<m}C_{S}(\Delta)$. (The constants depend on $p,q,T,d,\underline{\nu},\|a^{*}\|_{\infty},\|b^{*}\|_{\infty}$ only through $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$.)
@@
From Lemma 3, Lemma 2, and Proposition C.1, we have the following basic three inequalities:
$E_{f}$ $\;\leq\;C_{f}\,\varepsilon+C_{f}^{\prime}E_{v},$ (36a)
$E_{h}$ $\;\leq\;C_{\mathrm{FOC}}\varepsilon+L_{H}E_{v},$ (36b)
$E_{v}$ $\;\leq\;C_{S}\,(E_{f}+E_{h}).$ (36c)
Inserting (36a) and (36b) into (36c) yields a single scalar inequality for $E_{v}$:
$E_{v}\;\leq\;C_{S}\Bigl{[}(C_{f}+C_{\mathrm{FOC}})\varepsilon+(C_{f}^{\prime}+L_{H})E_{v}\Bigr{]}\;=\;C_{S}(C_{f}+C_{\mathrm{FOC}})\varepsilon+\rho\,E_{v}.$
Rearranging this gives $(1-\rho)\,E_{v}\;\leq\;C_{S}(C_{f}+C_{\mathrm{FOC}})\,\varepsilon$.
Now, we consider two cases for $\rho$. If $\rho<1$, the inequality immediately gives $E_{v}\leq\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}\,\varepsilon$. Define $C_{v}^{\prime\prime}:=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}$. Back‐substituting this into (36a) and (36b) yields (35) with the constants stated for the $\rho<1$ branch, setting $\rho_{\!*}:=\rho$.
If $\rho\geq 1$, we use a time–slicing argument. Fix $m\in\mathbb{N}$ and set $\Delta:=T/m$. Let $E_{v}^{(k)}$ denote the $L^{q,p}$–norm of $\nabla V^{\pi,c}-\nabla V^{*}$ on the slab $Q_{k}:=(T-k\Delta,\,T-(k-1)\Delta)\times\Omega$. Because the coefficients are effectively frozen on each slab, Proposition C.1 holds with a short‐horizon constant $C_{S}(\Delta)$, which satisfies $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$ (as the Dong–Kim estimate is local in time). Repeating the derivation for the scalar inequality for $E_{v}$ on $Q_{k}$ yields $E_{v}^{(k)}\;\leq\;\rho_{\Delta}\,E_{v}^{(k-1)}+C_{S}(\Delta)\bigl{(}C_{f}+C_{\mathrm{FOC}}\bigr{)}\varepsilon$, where $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$. We pick $m$ such that $\rho_{\Delta}<1$. Iterating for $k=1,\dots,m$ and noting $E_{v}^{(0)}=0$ (due to the terminal condition $W(T,\cdot)=0$), we get $E_{v}^{(m)}\;\leq\;\frac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}\varepsilon$. With $C_{v}^{\prime\prime}:=\dfrac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}$, $\rho_{\!*}:=\rho_{\Delta}$, and $C_{S}$ redefined as $\max_{0\leq k<m}C_{S}(\Delta)$ for the overall bound, the results (35) follow similarly to the $\rho<1$ case.
In both scenarios, all quantities $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}(\Delta)$ depend only on the data in Assumption 6. In particular, $C_{S}(\Delta)\leq C_{S}(T)$, so $C_{v}^{\prime\prime},C_{f}^{\prime\prime},C_{h}^{\prime\prime}$ are finite. ∎
Theorem 4 (Policy Gap Bound).
Under Assumptions 6 and 7, let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$. Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs. Assume $F$ is locally Lipschitz continuous with constant $C_{F}$ (guaranteed under Assumption 6(c)). If the BPTT estimation error relative to the Stage 1 costate satisfies $\|\widehat{\lambda}-\lambda\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$, then the policy gap is bounded by:
$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
Here, $C_{\mathrm{tot}}$ depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ from Proposition C.2.
Proof (detailed).
Let $(\lambda_{S1},Z_{S1})$ denote the adapted solution (specifically, $\lambda_{S1}=\nabla V^{\pi,c}$) associated with the Stage 1 policy $(\pi,c)$. The BPTT process yields estimates $\widehat{\lambda}$ (and implicitly $\widehat{Z}$) for the costate and its related processes. The first–order map $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$ is $C^{1}$ on a neighbourhood $\mathcal{N}$ of $(\lambda^{*},\nabla\lambda^{*})$ with Lipschitz constant $\|F\|_{\mathrm{Lip}}\leq C_{F}:=\underline{\Lambda}^{-1}L_{HX}$ (derived similarly to reasoning in Lemma 2, under Assumption 6(c)).
The policy deployed in Stage 2 is $\widehat{\pi}=F(\widehat{\lambda})$, and the true optimal policy is $\pi^{*}=F(\nabla V^{*})$. We want to bound $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$. Using the Lipschitz continuity of $F$:
$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$ $=\|F(\widehat{\lambda})-F(\nabla V^{*})\|_{L^{q,p}}$
$\leq C_{F}\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}.$
By the triangle inequality,
$\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}\leq\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}+\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}.$
We are given that the BPTT estimation error $\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$. The term $\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}$ is precisely $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$. From Proposition C.2, we have $E_{v}\leq C_{v}^{\prime\prime}\varepsilon$. Substituting these into the inequality for $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$:
@@
$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\leq C_{F}\left(\delta_{\mathrm{BPTT}}+E_{v}\right)\leq C_{F}\left(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon\right).$
To express this in the form $C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)$, we can set $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$. If $C_{v}^{\prime\prime}\geq 1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}C_{v}^{\prime\prime}(\delta_{\mathrm{BPTT}}+\varepsilon)$. If $C_{v}^{\prime\prime}<1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(\delta_{\mathrm{BPTT}}+\varepsilon)$. Thus, $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ provides a valid constant for the boxed expression. Alternatively, one can use $C_{\mathrm{tot}}=C_{F}(1+C_{v}^{\prime\prime})$ if one prefers a sum form, as $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(1+C_{v}^{\prime\prime})(\delta_{\mathrm{BPTT}}+\varepsilon)$ assuming $C_{v}^{\prime\prime}\geq 0$. The choice $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ is tighter. The key result is the linear dependence on $\delta_{\mathrm{BPTT}}$ and $\varepsilon$. ∎
4. Summary and Constant Dependencies.
We have demonstrated a clear logical path from a small FOC-gap $\varepsilon$ to a small policy gap $O(\varepsilon+\delta_{\mathrm{BPTT}})$, conditional on standard parabolic regularity. The argument relies on establishing key bounds (Lemmas 2, 3), utilizing PDE stability (Proposition C.1), resolving coupled estimates (Proposition C.2), and concluding with the main policy gap theorem (Theorem 4).
It is important to reiterate that all constants ($L_{H},C_{\mathrm{FOC}},C_{f},C_{f}^{\prime},C_{S},$ $C_{f}^{\prime\prime},C_{h}^{\prime\prime},C_{v}^{\prime\prime},C_{F},C_{\mathrm{tot}}$) depend quantitatively on model parameters (horizon $T$, utility $U$, market coefficients $\bm{\mu},\bm{\sigma},\dots$), PDE regularity parameters ($p,q,\Omega,\underline{\nu}$), and the costate floor $\underline{\lambda}$. Rigorous derivation requires tracking these dependencies. The provided proofs for lemmas are sketches; full proofs require careful application of MVT, IFT, Taylor expansions, and meticulous handling of bounds. The time-slicing argument mentioned requires careful justification to control constant growth. Subject to these standard technical details, the framework presented provides a sound mathematical basis for the Projected PG-DPO method’s effectiveness under the stated regularity conditions.

### v4

@@ C.4 Automatic Properties of the Model
Apply Dong and Kim, (2009) to the non‑divergence PDE for $V^{\pi,c}$ using the VMO modulus bound, uniform ellipticity and bounded coefficients. The same argument holds for $V^{*}$. ∎
C.5 Foundational Lemmata
Lemma 6 (FOC‑Reduction: Coefficient Gap).
Under Assumption 6 and Lemma 4,
$E_{h}\;\leq\;L_{H}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
@@ C.5 Foundational Lemmata
The strong concavity of the Hamiltonian (Lemma 4) implies that its gradient with respect to the control, $G(u;\lambda):=\nabla_{u}H(u,\lambda)$, is a strongly monotone operator in $u$ for a fixed state $\lambda$. By the Minty–Browder theorem (Minty, , 1962; Browder, , 1965), such an operator has a Lipschitz-continuous inverse. This allows us to bound the gap in controls by the gap in costates ($E_{v}$) and the FOC residual ($\varepsilon$), which after applying the control-to-coefficient Lipschitz bound, yields the stated inequality. ∎
Lemma 7 (Source‑Gap).
Under Assumption 6,
$E_{f}\;\leq\;C_{f}^{\prime}E_{v}+C_{f}\varepsilon.$
@@
Compactness of the costate band and the properties of HARA utility imply that both $U$ and $(U^{\prime})^{-1}$ are Lipschitz continuous on the relevant compact intervals. The result follows from a standard mean-value argument on the first-order conditions for consumption. ∎
Lemma 8 (Gradient Stability).
With Lemma 5 in force,
$E_{v}\;\leq\;C_{S}(E_{h}+E_{f}).$
Proof.
Let $W:=V^{\pi,c}-V^{*}$. The function $W$ solves a linear parabolic PDE whose source term is determined by the coefficient and source gaps ($h$ and $f$). By the Dong–Kim gradient estimate for PDEs with VMO coefficients, a core result of maximal regularity theory, the gradient of the solution is bounded by the norm of the source term, which gives $\|\nabla W\|_{L^{q,p}}\leq C_{S}(E_{h}+E_{f})$. ∎
Proposition 1 (Coupled Error Resolution).
There exists $C_{v}^{\prime\prime}$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
Proof.
Combining the three preceding lemmas yields the recursive inequality $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, where $\alpha=C_{S}(L_{H}+C_{f}^{\prime})$. For a sufficiently small time interval $\Delta t\leq\Delta_{\star}$, PDE theory ensures that the constant $\alpha<1$. This turns the mapping $T(E_{v})=\alpha E_{v}+\beta\varepsilon$ into a contraction. By the contraction mapping principle, this recursion converges. Iterating this argument over the $\lceil T/\Delta_{\star}\rceil$ time slabs that cover the full horizon $[0,T]$ yields a final bound for $E_{v}$ that depends only on $\varepsilon$. ∎
C.6 Final Policy‑Gap Theorem
Definition 1 (BPTT Error).
$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\kappa_{1}\Delta t+\kappa_{2}/\sqrt{M},$ with constants depending on the value function regularity established in Lemma 5.
Theorem 1 (Policy Gap Bound). Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$. Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
@@ C.6 Final Policy‑Gap Theorem
$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
where:
  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
Proof.
The proof follows from the triangle inequality: $\|\widehat{\pi}-\pi^{*}\|\leq L_{F}\bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{*}\|\bigr{)}\leq L_{F}(\delta_{\mathrm{BPTT}}+E_{v}).$ Substituting the bounds for $\delta_{\mathrm{BPTT}}$ from Definition 1 and for $E_{v}$ from Proposition 1 yields the final result. ∎